I am the worst woth word problems. Could someone help me with this one.
During rush hour, Fernando can drive 35 miles using the side roads in the same time that it takes to travel 25 miles on the freeway. If Fernando's rate on the side roads is 8 mi/h faster than his rate on the freeway, find his rate on the side roads.

Answer by scott8148(6628)

let s=(rate on side roads), so (25/35)s=(rate on freeway)

s=(25/35)s+8 ... 35s=25s+280 ... 10s=280 ... s=28